# VAMO: Efficient Large-Scale Nonconvex Optimization via Adaptive Zeroth Order Variance Reduction

### v1

@@ A.6 Auxiliary Lemmas
This is a direct consequence of Lemma A.2 in lei2017non . ∎
A.7 Analysis of Zeroth-Order Gradient Estimation Error
This section details bounds on the expected squared error of the ZO gradient estimators used in our work. We consider a ZO gradient estimator ^$\hat{\nabla}f_{i}(\mathbf{x})$ for a component function $f_{i}(\mathbf{x})$, which approximates the true gradient $\nabla f_{i}(\mathbf{x})$ with an estimation error $\omega_{i}(\mathbf{x})$, such that ^$\hat{\nabla}f_{i}(\mathbf{x})=\nabla f_{i}(\mathbf{x})+\omega_{i}(\mathbf{x})$. The characteristics of the expected squared error, $\mathbb{E}[\|\omega_{i}(\mathbf{x})\|_{2}^{2}]$, are presented below.
For the two-point ZO gradient estimator of $f_{i}(\mathbf{x})$, as defined in Equation (2) in the main text, the expected squared error is bounded by:
@@ A.7 Analysis of Zeroth-Order Gradient Estimation Error
Subsequently, for the multi-point ZO gradient estimator of $f_{i}(\mathbf{x})$ using $2q$ query points, as defined in Equation (3) in the main text, the expected squared error is bounded by:
@@
The detailed proofs for these bounds can be found in Proposition 2 of liu2018zeroth .
A.8 Nonconvex Least Squares Task
@@
The primary objective of this experiment was to empirically investigate the impact of the number of ZO query points ($q$) on the performance of VAMO and to validate the theoretical benefits of its multi-point ZO estimation strategy. The optimization problem was a finite-sum non-convex least-squares objective: $f(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}(h(\mathbf{x};\mathbf{z}_{i})-y_{i})^{2}$. We configured this synthetic task with $n=1000$ individual component functions and a parameter dimension of $d=100$. The function $h(\mathbf{x};\cdot)$ was parameterized using a simple neural network with a non-convex activation function to ensure the overall non-convexity of the loss landscape.
@@
In this setup, VAMO variants utilizing $q\in\{1,3,5\}$ query directions for the multi-point ZO gradient estimator were compared against the classical first-order SGD algorithm. A mini-batch size of $b=8$ was consistently applied across all methods. Learning rates for both VAMO (for each $q$ setting) and FO-SGD were individually tuned by selecting the best performing value from the range $[10^{-2},10^{-1}]$. For all VAMO variants, the ZO smoothing parameter was fixed at $\mu=10^{-3}$. The mixing coefficient $\alpha$ for VAMO was also tuned for each value of $q$, guided by the theoretical insights on balancing FO and ZO information discussed in Section 5.1.
A.9 MNIST Classification Task
@@
For the MNIST multi-class image classification task lecun1998gradient , we trained a Multi-Layer Perceptron (MLP) to evaluate VAMO against established baselines. The MLP architecture consisted of an input layer receiving flattened $28\times 28$ pixel images (784 dimensions), followed by two hidden layers with 32 and 16 units respectively, both employing ReLU activation functions. The final output layer comprised 10 units corresponding to the digit classes, and the network was trained using a standard cross-entropy loss function. Images were normalized to the range $[0,1]$.
@@
Our proposed VAMO algorithm, configured with a single ZO query direction ($q=1$), was benchmarked against pure first-order (FO-SGD) robbins1951stochastic and pure zeroth-order methods (ZO-SGD and ZO-SVRG) liu2018zeroth ; ghadimi2013stochastic . For all methods, the mini-batch size was set to $b=4$. Learning rates were independently tuned for each method, selected from the range $[10^{-4},10^{-3}]$ for ZO methods, and $[10^{-3},10^{-2}]$ for FO methods and VAMO. For VAMO with $q=1$, we fixed the mixing coefficient at $\alpha=0.1$ and used a ZO smoothing parameter of $\mu=10^{-3}$. All models were trained for 10 epochs.
A.10 GPT2 Fine-Tuning on MNLI
@@
Experiment Setup This experiment was designed to evaluate VAMO’s performance in the practical and challenging context of fine-tuning large language models. We fine-tuned a pre-trained GPT-2 model radford2019language on the MultiNLI (MNLI) dataset williams2017broad for a three-way natural language inference task. The training set was subsampled to 256 examples and the validation set to 128 examples, with a maximum input sequence length of 512 tokens. All models were fine-tuned for 1000 epochs, using half-precision (FP16) training for computational efficiency. All experiments were conducted on a single NVIDIA A100 GPU with 40 GB memory.
VAMO was benchmarked against several representative FO and ZO methods. To ensure a fair comparison, we tuned the learning rates for FO methods in the range $[10^{-4},10^{-3}]$, and for ZO methods in the range $[10^{-6},10^{-5}]$, based on their respective convergence behaviors. For the FO-SGD method robbins1951stochastic , we used an effective batch size of 32. For ZO methods such as ZO-SGD ghadimi2013stochastic and ZO-SVRG liu2018zeroth , we adopted a smoothing parameter $\mu=10^{-3}$ and $q=1$ (two-point version) query direction per iteration. In the primary comparison, VAMO was configured with a batch size of 32, smoothing parameter $\mu=10^{-3}$, and $q=1$. The mixing coefficient was set to $\alpha=0.05$, based on the analysis in Section 5.1. The number of inner loop iterations was set to $m=10$. No learning rate scheduler was applied to VAMO to isolate the effect of its variance reduction mechanism.
Role of $m$ (Inner Loop Iterations): To study the impact of inner loop length $m$, we fixed $\alpha=0.1$, $q=1$, and $\mu=10^{-3}$, and varied $m\in\{2,5,10\}$. These correspond to different frequencies of ZO checkpointing. We report the final validation accuracy after 1000 epochs to evaluate the trade-off between variance reduction effectiveness and ZO query overhead. Results are shown in Figure 3.
Role of $\alpha$ (Mixing Coefficient): We varied $\alpha\in\{0.05,0.1,0.15,0.2\}$ to investigate how strongly ZO information should be incorporated at SVRG checkpoints. During these runs, other VAMO parameters were fixed. As shown in Figure 3, under our experimental settings, we found that a moderate value of $\alpha=0.05$ achieved a favorable trade-off between fast convergence and robustness to ZO estimation noise, consistent with the analysis presented in Section 5.1.

### v2

@@ G.1 Auxiliary Lemmas
This is a direct consequence of Lemma A.2 in (Lei et al., 2017). ∎
Appendix H Analysis of Zeroth-Order Gradient Estimation Error
This section details bounds on the expected squared error of the ZO gradient estimators used in our work. We consider a ZO gradient estimator $\hat{\nabla}f_{i}(\mathbf{x})$ for a component function $f_{i}(\mathbf{x})$, which approximates the true gradient $\nabla f_{i}(\mathbf{x})$ with an estimation error $\omega_{i}(\mathbf{x})$, such that $\hat{\nabla}f_{i}(\mathbf{x})=\nabla f_{i}(\mathbf{x})+\omega_{i}(\mathbf{x})$. The characteristics of the expected squared error, $\mathbb{E}[\|\omega_{i}(\mathbf{x})\|_{2}^{2}]$, are presented below.
For the two-point ZO gradient estimator of $f_{i}(\mathbf{x})$, as defined in Equation equation 2 in the main text, the expected squared error is bounded by:
@@ Appendix H Analysis of Zeroth-Order Gradient Estimation Error
Subsequently, for the multi-point ZO gradient estimator of $f_{i}(\mathbf{x})$ using $2q$ query points, as defined in Equation equation 3 in the main text, the expected squared error is bounded by:
@@
The detailed proofs for these bounds can be found in Proposition 2 of (Liu et al., 2018b).
Appendix I Experiment Setup
I.1 Adaptability Experiment
@@
The primary objective of this adaptability experiment was to empirically investigate the impact of the number of ZO query points ($q$) on the performance of VAMO and to validate the theoretical benefits of its multi-point ZO estimation strategy (see Section 5). The optimization problem was a finite-sum non-convex least-squares objective: $f(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}(h(\mathbf{x};\mathbf{z}_{i})-y_{i})^{2}$. We configured this synthetic task with $n=1000$ individual component functions and a parameter dimension of $d=100$. The function $h(\mathbf{x};\cdot)$ was parameterized using a simple neural network with a non-convex activation function to ensure the overall non-convexity of the loss landscape.
@@
In this setup, VAMO variants utilizing $q\in\{1,3,5\}$ query directions for the multi-point ZO gradient estimator were compared against the classical FO-SGD algorithm. A mini-batch size of $b=8$ was consistently applied across all methods. Learning rates for both VAMO (for each $q$ setting) and FO-SGD were individually tuned by selecting the best performing value from the range $10^{-2},10^{-1}]$. For all VAMO variants, the ZO smoothing parameter was fixed at $\mu=10^{-3}$. The mixing coefficient $\alpha$ for VAMO was also tuned for each value of $q$, guided by the theoretical insights on balancing FO and ZO information discussed in Appendix J.2.
I.2 MNIST Classification Task
@@
For the MNIST multi-class image classification task (LeCun et al., 1998), we trained a Multi-Layer Perceptron (MLP) to evaluate VAMO against established baselines. The MLP architecture consisted of an input layer receiving flattened $28\times 28$ pixel images (784 dimensions), followed by two hidden layers with 32 and 16 units respectively, both employing ReLU activation functions. The final output layer comprised 10 units corresponding to the digit classes, and the network was trained using a standard cross-entropy loss function. Images were normalized to the range $[0,1]$.
@@
Our proposed VAMO algorithm, configured with a single ZO query direction ($q=1$), was benchmarked against pure first-order FO-SGD (Robbins & Monro, 1951) and pure zeroth-order methods, ZO-SGD and ZO-SVRG  (Liu et al., 2018b; Ghadimi & Lan, 2013). For all methods, the mini-batch size was set to $b=4$. The learning rates were independently tuned for each method, selected from the range $10^{-4},10^{-3}]$ for the ZO methods, and $10^{-3},10^{-2}]$ for FO methods and VAMO. For VAMO with $q=1$, we fixed the mixing coefficient at $\alpha=0.1$ and used a ZO smoothing parameter of $\mu=10^{-3}$. All models were trained for 10 epochs.
I.3 Fine-Tuning Experiments
@@
To further examine VAMO in realistic large-scale scenarios, we conducted fine-tuning experiments on three language models under the MultiNLI (MNLI) dataset (Williams et al., 2017) for a three-way natural language inference task. In all settings, the training and validation sets were subsampled to 256 and 128 examples, respectively, with a maximum input sequence length of 128 tokens. All models were fine-tuned for 1500 epochs with full precision (FP32) on a single NVIDIA RTX5880 49GB GPU. All experiments involved full-parameter fine-tuning.
For fair comparisons, VAMO was compared against representative FO methods including FO-SGD (Robbins & Monro, 1951), FO-Adagrad (Duchi et al., 2011) and FO-Adam (Kingma & Ba, 2014), and ZO methods including ZO-SGD (Ghadimi & Lan, 2013) and ZO-SVRG (Liu et al., 2018b). We set the batch size as 32. The learning rates of FO methods were tuned within $10^{-4},10^{-3}]$, while those of ZO methods were tuned within $10^{-6},10^{-5}]$. For VAMO, the learning rate was also tuned in the range $10^{-4},10^{-3}]$. We adopted the same learning rate schedule as illustrated in Algorithm 2. For ZO methods, we adopted a smoothing parameter $\mu=10^{-3}$ and $q=1$ (two-point version). VAMO was configured with the the smoothing parameter $\mu=10^{-3}$, $q=1$, and inner loop length $m=10$.
The main difference is in setting the mixing coefficient $\alpha$. $\alpha$ was set to $0.01$ for the GPT-2 experiments, while a lower value $\alpha=0.005$ was used for RoBERTa-Large and GPT-2 Medium. We used a smaller $\alpha$ for larger models because the ZO gradient estimates exhibit higher inherent error in higher-dimensional models, so reducing $\alpha$ limits the contribution of the noisier ZO correction and helps stabilize the training. A detailed analysis of $\alpha$ selection and the error of the ZO estimation is provided in Appendix J.2 and Appendix H.
Algorithm 2 Learning Rate Scheduling for VAMO
1:Input: Learning rates $\eta_{1},\eta_{2}$, annealing factor $\delta_{\eta}$, losses $L$, annealing threshold $\kappa$, total number of batches in an epoch $w$
2: Compute moving averages:
[…]
5:if $\frac{L_{1}}{L_{2}}>\kappa$ then
6:  $\eta_{1}\leftarrow\frac{\eta_{1}}{\delta_{\eta}},\quad\eta_{2}\leftarrow\frac{\eta_{2}}{\delta_{\eta}}$
7:end if
8:Return: updated $\eta_{1},\eta_{2}$
Appendix J Ablation on Key Hyperparameters
In this section, we investigate the effect of two critical hyperparameters in VAMO: the number of inner loop iterations $m$ and the mixing coefficient $\alpha$. Both parameters play a central role in balancing the trade-offs of the algorithm—$m$ controls the frequency of full-batch ZO gradient corrections, while $\alpha$ adjusts the relative contribution of ZO-based variance reduction. To better understand their influence, we conduct ablation studies on fine-tuning GPT-2 Medium (Radford et al., 2019) under the MNLI dataset (Williams et al., 2017), examining both convergence speed and stability.
J.1 Role of $m$ (Inner Loop Iterations)
The parameter $m$ is a significant role in VAMO as it governs the frequency of the fullbatch ZO gradient updates. Specifically, a smaller $m$ means that the full-batch ZO gradient estimate is computed more frequently, which can lead to a more effective reduction in gradient variance. To better understand the trade-off between the effectiveness of these updates and their computational overhead, we perform an ablation study on $m$.
We consider the task of fine-tuning a GPT-2 model on the MNLI dataset . For this study, we fixed the other key hyperparameters ($\alpha=0.1$, $q=1$, and $\mu=10^{-3}$) while varying $m$. The results, illustrated in Figure 3, show that increasing the frequency of fullbatch ZO gradient update ($2\leq m\leq 10$) generally enhances the performance of VAMO. However, this trend does not continue indefinitely; when a full batch update becomes too infrequent (e.g., $m\geq 20$), the variance correction becomes stale and less effective, leading to a performance degradation. This suggests an optimal range for $m$ that balances the benefits of frequent variance correction against its computational cost.
J.2 Role of $\alpha$ (Mixing Coefficient)
Theoretical Analysis.
[…]
Empirical Results Analysis.
The mixing coefficient $\alpha$ controls how much weight VAMO assigns to the ZO-based variance reduction term relative to the FO gradient. A larger $\alpha$ strengthens the effect of variance correction but also amplifies the noise of the ZO estimator. To study this trade-off, we fine-tuned GPT-2 on the MNLI dataset, fixing other hyperparameters ($m=10$, $q=1$, $\mu=10^{-3}$) while varying $\alpha$. As shown in Figure 3, smaller values (e.g. $\alpha=0.01$) provide a good balance, leading to fast convergence without being overly sensitive to noisy ZO gradients. However, if $\alpha$ is set too small (e.g., $\alpha=0.005$), the variance reduction term becomes underutilized, diminishing its benefit. In contrast, setting $\alpha$ too high (e.g., $\alpha=0.2$) magnifies estimation error, destabilizes training, and slows convergence. These results align with our theoretical analysis, which highlights the need for careful tuning of $\alpha$ in high-dimensional settings.